OpenStudy (anonymous):

don has 18 coins. some are quarters and some are dimes. the total value of the coins is $3.45. How many of each type of coin does he have? Be sure to identify your variables

OpenStudy (anonymous):

\[25x+(10(18-x)=345\] solve for x

OpenStudy (anonymous):

typo on first parentheses

OpenStudy (watchmath):

Let \(x,y\) be the number of quarters and dimes respectively. Then \(x+y=18\) and \(25x+10y=345\)(cents) Mltiply the first by -25 and add it to the second equation we have \(-15y=-105\). So \(y=7\). Plug in to the first equation \(x+7=11\). So \(x=11\)

OpenStudy (anonymous):

wouldn't i do .25x+.10y=3.45

OpenStudy (watchmath):

typo \(x+7=18\) so \(x=11\)

OpenStudy (anonymous):

i would not use two variables when one can do, but this is a matter of preference

OpenStudy (watchmath):

what is this blue little circle means satellite?

OpenStudy (anonymous):

dunno this is all new to me

OpenStudy (anonymous):

i think they should give amistre his own special hypersphere

OpenStudy (anonymous):

i feel BAD FOR U POLLOCK DONT HELP HES JUST MEAN